OpenStudy (aundiewear):

Use an appropriate Half-Angle Formula to find the exact value of the expression. sin(13pi/12)

OpenStudy (dumbcow):

\[\sin(\frac{13\pi}{12}) = -\sin(\frac{\pi}{12}) = -\sqrt{\frac{1-\cos(\frac{\pi}{6})}{2}}\]